INTERACTION BETWEEN VIBRATIONS AND INTERNAL ROTATION IN S_{1} M-FLUOROTOLUENE AS SEEN IN THE S_{1}-S_{0} SPECTROSCOPY

\underline{QUAN JU}, AND CHARLES S. PARMENTER, Department of Chemistry, Indiana University, Bloomington, IN 47405.

In support of IVR studies in m-fluorotoluene, a detailed picture is being constructed of the vibrational and internal rotation level structure for the lower 430 cm^{−1} of the S1 state. Particular emphasis is given to Fermi resonances involving interactions between internal rotation and vibration. The information is derived from cold jet S_{1}-S_{0} spectroscopy with abundant use of dispersed single vibronic level fluorescence (SVLF). Seventy-five levels involving internal rotation as well as eight modes occurring as fundamentals, overtones and combinations occur with known zero order identities in the energy region. Of these, twenty-six have been pumped for examination by SVLF spectroscopy. Clear evidence for numerous level interactions appears in the spectra. The details are now being worked out.